OpenStudy (anonymous):

what is the equation in standard form of a perpendicular line that passes through 5,1? y+4=-4(x-2)

OpenStudy (lucaz):

you need a slope, the perpendicular line has the negative reciprocal slope

OpenStudy (anonymous):

i keep getting the answer wrong

OpenStudy (lucaz):

what is the slope of the line that passes through the point (5,1) ?

OpenStudy (lucaz):

you need this information

OpenStudy (anonymous):

2/-4

OpenStudy (anonymous):

right?

OpenStudy (lucaz):

\[\large{ -\frac{2}{4} }\] this is the slope?

OpenStudy (anonymous):

yes im pretty sure

OpenStudy (lucaz):

it simplifies to -1/2, so now you get the reciprocal neg. slope -1/(-1/2) = 2 and use the standard form \[\large{ y-y_1=m(x-x_1)}\]

OpenStudy (lucaz):

this y+4=-4(x-2) is the given line?

OpenStudy (anonymous):

yes that is what i did and got y-1=2/-4(x-5)

OpenStudy (lucaz):

if this is the given line the slope is -4, and it's neg. reciprocal is 1/4

OpenStudy (anonymous):

but i dont get what y equals? because in another question i have to graph the points?

OpenStudy (anonymous):

and i got x=1 but i dont know what y is?

OpenStudy (lucaz):

this is just a function, if you have x=1 you replace x by 1 and calculate y

OpenStudy (anonymous):

but what about 5,1?

OpenStudy (lucaz):

like f(x) or y = 3x+1, if x=1 -> 3(1)+1 = 4

OpenStudy (lucaz):

yeah, the slope is -4, so the reciprocal is 1/4